# A Non-Commutative Kalb-Ramond Black Hole

### v2

@@ Abstract
This work presents a new black hole solution within the framework of a non–commutative gauge theory applied to Kalb–Ramond gravity. Using the method recently proposed in the literature [Nucl.Phys.B 1017 (2025) 116950], we employ the Moyal twist $\partial_{r}\wedge\partial_{\theta}$ to implement non–commutativity, being encoded by parameter $\Theta$. We begin by verifying that the resulting black hole no longer possesses spherical symmetry, while the event horizon remains unaffected by non–commutative corrections. The Kretschmann scalar is computed to assess the corresponding regularity. It turns out that the solution is regular, provided that the Christoffel symbols and related quantities are not expanded to second order in $\Theta$. We derive the thermodynamic quantities, including the Hawking temperature $T^{(\Theta,\ell)}$, entropy $S^{(\Theta,\ell)}$, and heat capacity $C_{V}^{(\Theta,\ell)}$. The remnant mass $M_{\text{rem}}$ is estimated by imposing $T^{(\Theta,\ell)}\to 0$, although the absence of a physical remnant indicates complete evaporation. Quantum radiation for bosons and fermions is analyzed via the tunneling method, where divergent integrals are treated using the residue theorem. Notably, in the low–frequency regime, the particle number density for bosons surpasses that of fermions (at least within the scope of the methods considered here). The effective potential for a massless scalar field is obtained perturbatively, enabling the computation of quasinormal modes and the time–domain profiles. Finally, further bounds on $\Theta$ and $\ell$ (Lorentz–violating paramter) are derived from solar system tests, including the perihelion precession of Mercury, light deflection, and the Shapiro time delay.
@@ I Introduction
Gravitational phenomena in general relativity are described through the curvature of spacetime, where the nonlinearity of Einstein’s field equations poses substantial difficulties for obtaining analytic solutions. Even when symmetries or simplifying assumptions are introduced, the task of solving these equations exactly remains largely intractable misner1973gravitation ; wald2010general . A common strategy to address these challenges involves linearizing the equations under the assumption of a weak gravitational field. This perturbative approach facilitates the analysis of spacetime fluctuations, particularly gravitational waves. These perturbations are crucial in the context of black hole dynamics, as they affect processes such as Hawking radiation, the system’s stability, and its interaction with the surrounding medium.
@@
The concept of a non–commutative spacetime stems from redefining the fundamental relations between coordinates, replacing their classical behavior with the commutation rule $[x^{\mu},x^{\nu}]=\mathbbm{i}\Theta^{\mu\nu}$, where the antisymmetric matrix $\Theta^{\mu\nu}$ quantifies the deformation. This shift away from traditional geometrical assumptions has led to several methods for embedding non–commutativity within gravitational theories. Among these, a notable route consists in modifying the underlying gauge structure—extending the SO(4,1) de Sitter symmetry and blending it with the Poincaré group ISO(3,1)—by employing the Seiberg–Witten map. Within this context, Chaichian and collaborators chaichian2008corrections introduced a version of the Schwarzschild solution that incorporates non–commutative corrections to the spacetime geometry.
@@
$-\frac{1}{16}\Theta^{\nu\rho}\Theta^{\lambda\tau}\Big{[}\tilde{\omega}^{ac}_{\nu}\tilde{\omega}^{cd}_{\lambda}\Big{(}D_{\tau}R^{d5}_{\rho\mu}+\partial_{\tau}R^{d5}_{\rho\mu}\Big{)}\Big{]}.$
@@
Two distinct Kalb–Ramond black hole configurations have emerged in recent theoretical investigations. The first, originally introduced in Yang:2023wtu , has become the subject of a broad array of studies encompassing multiple physical aspects. Among them are the analysis of quasinormal spectra araujo2024exploring , lensing signatures in strong field regimes junior2024gravitational and greybody emissions guo2024quasinormal . Researchers have also examined how this geometry accommodated spontaneous symmetry–breaking effects junior2024spontaneous (by addressing bounds), kinetic theory scenarios involving Vlasov gas accretion jiang2024accretion and particle dynamics in circular orbits and quasi–periodic oscillations jumaniyozov2024circular . Investigations have further extended to mechanisms of particle production 12araujo2024particle , global monopole structures belchior2025global , and slowly rotating extensions of the solution Liu:2024lve . Building upon this framework, an electrically charged version has been constructed duan2024electrically , and additional features concerning thermal aspects, lensing phenomena, radiation profiles, and topological remarks were also taken into account al-Badawi:2024pdx ; Pantig:2025eda ; heidari2024impact ; aa2024antisymmetric ; chen2024thermal ; Zahid:2024ohn ; hosseinifar2024shadows . A second solution, proposed more recently in Liu:2024oas , has initiated interest in the degradation of entanglement under Lorentz symmetry violation liu2024lorentz , as well as in quantum field processes such as particle emission and greybody corrections 12araujo2024particle ; 12araujo2025does .
@@
This paper is organized as follows: In Sec. II, we present the general framework of the non–commutative gauge theory and outline the steps for constructing a new black hole solution within Kalb–Ramond gravity. We also examine the effects of non–commutativity on the event horizon and the Kretschmann scalar. In Sec. III, we investigate the thermodynamic behavior, focusing on the Hawking temperature, entropy, and heat capacity, and estimate the remnant mass. In Sec. IV, we analyze the quantum radiation for bosons and fermions, where divergent integrals are handled using the residue method. In Sec. VI, we derive the effective potential for scalar perturbations using a perturbative approach. In Sec. VII, based on the effective potential, we apply the WKB method to compute the quasinormal modes. In Sec. VIII, we study the time evolution of perturbations by solving the time–domain profile. In Sec. V, we evaluate the evaporation lifetime as the black hole approaches its final stage, and compute the energy and particle emission rates in the high–frequency regime. In Sec. IX, we establish bounds on the parameters $\Theta$ and $\ell$ using solar system tests such as Mercury’s perihelion precession, light deflection, and the Shapiro time delay. Finally, in Sec. X, we present our concluding remarks.
@@ II The new non–commutative black hole
Initially, the analysis begins by outlining the underlying framework for treating gravity through the lens of a non–commutative gauge theory. The core of this construction lies the $\mathrm{SO}(4,1)$ group, associated with the de Sitter symmetry, which serves as the foundational gauge group. As a starting point, the formalism is developed in a standard $(3+1)$–dimensional Minkowski spacetime, still within a commutative setting. Within this background, the line element in spherical coordinates takes the familiar form:
@@
The metric’s angular sector is described by $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta,\mathrm{d}\varphi^{2}$, as one should expect. As mentioned before, the underlying symmetry is governed by the $\mathrm{SO}(4,1)$ group, whose algebra is generated by ten antisymmetric elements, labeled $\mathcal{M}_{ab}$ and satisfying the condition $\mathcal{M}_{ab}=-\mathcal{M}_{ba}$. Here, the indices $a$ and $b$ span the range ${0,1,2,3,5}$, with the subset ${0,1,2,3}$ denoted by lowercase Latin indices from the beginning of the alphabet. These generators can be separated into two distinct classes: $\mathcal{M}_{ab}$ (with $a,b=0,1,2,3$), associated with local Lorentz rotations, and $P_{a}=\mathcal{M}_{a5}$, which fundamentally are generators of spacetime translations.
@@
$e^{1}_{\mu}=\left(\frac{1}{\tilde{\mathcal{A}}},0,0,0\right),\quad e^{2}_{\mu}=\left(0,r,0,0\right),\quad e^{3}_{\mu}=\left(0,0,r\,\mathrm{sin}\theta,0\right),\quad e^{0}_{\mu}=\left(,0,0,0,\tilde{\mathcal{A}}\right),$ (4)
@@
$\begin{split}&T^{0}_{01}=-\frac{\tilde{\mathcal{A}}\tilde{\mathcal{A}}^{\prime}+\tilde{\mathcal{U}}}{\mathcal{A}},\qquad T^{2}_{03}=r\,\tilde{\mathcal{V}}\sin\theta\,T^{2}_{12}=\frac{\tilde{\mathcal{A}}+\tilde{\mathcal{W}}}{\tilde{\mathcal{A}}},\\ &T^{3}_{02}=-r\,\tilde{\mathcal{V}},\,\,\,\qquad\qquad T^{3}_{13}=\frac{\left(\tilde{\mathcal{A}}+\tilde{\mathcal{Z}}\right)\sin\theta}{\tilde{\mathcal{A}}}.\end{split}$ (6)
@@
Within this framework, the parameter $M$ identifies the mass of the black hole, while $\ell$ serves as the deformation parameter associated with the Kalb–Ramond sector Yang:2023wtu . The modified line element is written as $\mathrm{d}s^{2}=g^{(\Theta,\ell)}_{\mu\nu}(x,\Theta)\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}$, where the coordinates $x^{\mu}=(t,r,\theta,\varphi)$ describe a $(3+1)$–dimensional spacetime endowed with non–commutative structure. Constructing this geometry involves determining the deformed tetrad fields $\hat{e}^{a}_{\mu}(x,\Theta)$, which arise through a contraction mechanism linking the non–commutative gauge group $\mathrm{SO}(4,1)$ to the Poincaré group $\mathrm{ISO}(3,1)$, following the Seiberg–Witten map formalism 5 ; 4 ; 3 . The resulting spacetime, modified by non–commutativity, is subject to the following structural constraints:
@@
$\omega^{ab}_{\mu\nu\rho}(x)=\frac{1}{4}\left\{\tilde{\omega}_{\nu},\partial_{\rho}\tilde{\omega}{\mu}+R_{\rho\mu}\right\}^{ab},$ (13)
@@
[…]
@@
Here, the symbol $\ast$ stands for the conventional Moyal (star) product used in non–commutative geometry. Throughout the remainder of this analysis, we adopt natural units by setting $\hbar=c=G=1$. The corresponding line element can now be written explicitly as follows:
@@ III.1 Hawking temperature
Notice that by taking into account the temperature in terms of the mass is essential for probing whether a remnant mass ($M_{\text{rem}}$) emerges and for evaluating the black hole’s evaporation dynamics within the framework of the Stefan–Boltzmann law. While the detailed thermodynamic evolution—including aspects such as greybody factors, and absorption cross sections=—is of considerable interest, these topics extend beyond the objectives of the present work. It is important to emphasize that, as the effective potential will be constructed through a perturbative approach (as discussed later), determining these quantities—particularly the last one—is far from straightforward. The complexity involved, therefore, makes the calculation highly nontrivial.
@@
$M^{1}_{\text{rem}}=\frac{i\sqrt{3}\Theta}{(4-4\ell)\sqrt{\ell+1}},$ (35)
$M^{2}_{\text{rem}}=\frac{i\sqrt{3}\Theta}{4(\ell-1)\sqrt{\ell+1}}.$ (36)
@@ III.2 Heat capacity
$\begin{split}C^{(\Theta,\ell)}_{V}&=T\frac{\partial S}{\partial T}=-2\left(\pi r_{h}^{2}\right)-\frac{2\pi r^{2}\left(3\Theta^{2}+4(\ell+1)r_{h}^{2}\right)}{9\Theta^{2}+4(\ell+1)r_{h}^{2}}\\ &\approx\,-2\left(\pi r_{h}^{2}\right)+(3\pi-3\pi\ell)\Theta^{2}.\end{split}$ (37)
@@ IV.1 Bosonic particle modes
$\mathcal{S}=\int p_{\mu}\,\mathrm{d}x^{\mu}.$ (39)
When extracting the corresponding imaginary part, $\text{Im},\mathcal{S}$, it is the radial contribution that governs the result. The temporal term $p_{t},\mathrm{d}t=-\omega,\mathrm{d}t$ remains purely real and, therefore, does not contribute to the tunneling amplitude. As such, only the radial integration yields a nonzero imaginary component, which enable us to evaluate the tunneling probability as follows
@@
Let us define the Hamiltonian as $H=M-\omega^{\prime}$, with $\omega^{\prime}$ denoting the instantaneous energy of the particle being emitted. According to Hamilton’s equations, this leads to the differential relation $\mathrm{d}H=-\mathrm{d}\omega^{\prime}$. Integrating over the entire energy range of the emitted particle, from 00 to the total energy $\omega$, the imaginary part of the action becomes:
@@
[…]
@@
$\begin{split}&\text{Im}\,\mathcal{S}=\text{Im}\,\int_{0}^{\omega}-\mathrm{d}\omega^{\prime}\\ &\times\int_{r_{i}}^{r_{f}}\,\frac{\mathrm{d}r}{1-\frac{\sqrt{\frac{2\ell(M-\omega^{\prime})+\ell r-2(M-\omega^{\prime})}{\ell-1}}}{\sqrt{r}}+\frac{(4\ell(M-\omega^{\prime})^{2}-4(M-\omega^{\prime})^{2}+(M-\omega^{\prime})r)\Theta^{2}}{2\,r^{3}(2\ell(M-\omega^{\prime})-2(M-\omega^{\prime})+r)}+\Xi(\Theta,\ell,\omega^{\prime})},\end{split}^{\prime}$ (43)
@@
$\Gamma(\Theta,\ell,\omega)\sim e^{-2\,\text{Im}\,S}=e^{-4\pi(1-\ell)^{2}\omega(2M-\omega)-\frac{\pi(16\ell-17)\big{[}\ln M-\ln(M-\omega)\big{]}\Theta^{2}}{8(1-\ell)^{2}}}.$ (50)
@@
$n(\Theta,\ell,\omega)=\frac{\Gamma(\Theta,\ell,\omega)}{1-\Gamma(\Theta,\ell,\omega)}=\frac{1}{\exp\left\{4\pi(1-\ell)^{2}\omega(2M-\omega)+\frac{\pi(16\ell-17)\big{[}\ln M-\ln(M-\omega)\big{]}\Theta^{2}}{8(1-\ell)^{2}}\right\}-1}.$ (51)
Fig. 6 presents the distribution of bosonic particle production, denoted by $n_{\psi}$. As observed, both the non–commutative parameter $\Theta$ and the Lorentz–violating parameter $\ell$ increase the particle creation density. However, it is worth emphasizing that the system exhibits a stronger response to variations in $\ell$ than to changes in $\Theta$.

### v3

@@ Abstract
This work presents a new black hole solution within the framework of a non–commutative gauge theory applied to Kalb–Ramond gravity. Using the method recently proposed in the literature [Nucl.Phys.B 1017 (2025) 116950], we employ the Moyal twist $\partial_{r}\wedge\partial_{\theta}$ to implement non–commutativity, being encoded by parameter $\Theta$. We begin by verifying that the resulting black hole no longer possesses spherical symmetry, while the event horizon remains unaffected by non–commutative corrections. The Kretschmann scalar is computed to assess the corresponding regularity. It turns out that the solution is regular, provided that the Christoffel symbols and related quantities are not expanded to second order in $\Theta$. We derive the thermodynamic quantities, including the Hawking temperature $T^{(\Theta,\ell)}$, entropy $S^{(\Theta,\ell)}$, and heat capacity $C_{V}^{(\Theta,\ell)}$. The remnant mass $M_{\text{rem}}$ is estimated by imposing $T^{(\Theta,\ell)}\to 0$, although the absence of a physical remnant indicates complete evaporation. Quantum radiation for bosons and fermions is analyzed via the tunneling method, where divergent integrals are treated using the residue theorem. Notably, in the low–frequency regime, the particle number density for bosons surpasses that of fermions (at least within the scope of the methods considered here). The evaporation lifetime and emission rate are also examined. It turns out that non–commutativity accelerates the black hole evaporation process, making it faster compared to the Kalb–Ramond and Schwarzschild black holes. The effective potential for a massless scalar field is obtained perturbatively, enabling the computation of quasinormal modes and the time–domain profiles. Finally, further bounds on $\Theta$ and $\ell$ (Lorentz–violating paramter) are derived from solar system tests, including the perihelion precession of Mercury, light deflection, and the Shapiro time delay.
@@ I Introduction
Gravitational phenomena in general relativity are described through the curvature of spacetime, where the nonlinearity of Einstein’s field equations poses substantial difficulties for obtaining analytical solutions. Even when symmetries or simplifying assumptions are introduced, the task of solving these equations exactly remains largely intractable misner1973gravitation ; wald2010general . A common strategy to address these challenges involves linearizing the equations under the assumption of a weak gravitational field. This perturbative approach facilitates the analysis of spacetime fluctuations, particularly gravitational waves. These perturbations are crucial in the context of black hole dynamics, as they affect processes such as Hawking radiation, the system’s stability, and its interaction with the surrounding medium.
@@
The concept of a non–commutative spacetime stems from redefining the fundamental relations between coordinates, replacing their classical behavior with the commutation rule $x^{\mu},x^{\nu}]=\mathbbm{i}\Theta^{\mu\nu}$, where the antisymmetric matrix $\Theta^{\mu\nu}$ quantifies the deformation. This shift away from traditional geometrical assumptions has led to several methods for embedding non–commutativity within gravitational theories. Among these, a notable route consists in modifying the underlying gauge structure—extending the SO(4,1) de Sitter symmetry and blending it with the Poincaré group ISO(3,1)—by employing the Seiberg–Witten map. Within this context, Chaichian and collaborators chaichian2008corrections introduced a version of the Schwarzschild solution that incorporates non–commutative corrections to the spacetime geometry.
@@
$\frac{1}{16}\Theta^{\nu\rho}\Theta^{\lambda\tau}\Big{[}\tilde{\omega}^{ac}_{\nu}\tilde{\omega}^{cd}_{\lambda}\Big{(}D_{\tau}R^{d5}_{\rho\mu}+\partial_{\tau}R^{d5}_{\rho\mu}\Big{)}\Big{]}.$
@@
Two distinct Kalb–Ramond black hole configurations have emerged in recent theoretical investigations. The first, originally introduced in Yang:2023wtu , has become the subject of a broad array of studies encompassing multiple physical aspects. Among them are the analysis of quasinormal spectra araujo2024exploring , lensing signatures in strong field regimes junior2024gravitational , and greybody emissions guo2024quasinormal . Researchers have also examined how this geometry accommodated spontaneous symmetry–breaking effects junior2024spontaneous (by addressing bounds), kinetic theory scenarios involving Vlasov gas accretion jiang2024accretion and particle dynamics in circular orbits and quasi–periodic oscillations jumaniyozov2024circular . Investigations have further extended to mechanisms of particle production 12araujo2024particle , global monopole structures belchior2025global , and slowly rotating extensions of the solution Liu:2024lve . Building upon this framework, an electrically charged version has been constructed duan2024electrically , and additional features concerning thermal aspects, lensing phenomena, radiation profiles, and topological remarks were also taken into account al-Badawi:2024pdx ; Pantig:2025eda ; heidari2024impact ; aa2024antisymmetric ; chen2024thermal ; Zahid:2024ohn ; hosseinifar2024shadows . A second solution, proposed more recently in Liu:2024oas , has initiated interest in the degradation of entanglement under Lorentz symmetry violation liu2024lorentz , as well as in quantum field processes such as particle emission and greybody corrections 12araujo2024particle ; 12araujo2025does .
@@
This paper is organized as follows: In Sec. II, we present the general framework of the non–commutative gauge theory and outline the steps for constructing a new black hole solution within Kalb–Ramond gravity. We also examine the effects of non–commutativity on the event horizon and the Kretschmann scalar. In Sec. III, we investigate the thermodynamic behavior, focusing on the Hawking temperature, entropy, and heat capacity, and estimate the remnant mass. In Sec. IV, we analyze the quantum radiation for bosons and fermions, where divergent integrals are handled using the residue method. In Sec. V, we evaluate the evaporation lifetime as the black hole approaches its final stage, and compute the energy and particle emission rates in the high–frequency regime. In Sec. VI, we derive the effective potential for scalar perturbations using a perturbative approach. In Sec. VII, based on the effective potential, we apply the WKB method to compute the quasinormal modes. In Sec. VIII, we study the time evolution of perturbations by solving the time–domain profile. In Sec. IX, we establish bounds on the parameters $\Theta$ and $\ell$ using solar system tests such as Mercury’s perihelion precession, light deflection, and the Shapiro time delay. Finally, in Sec. X, we present our concluding remarks.
@@ II The new non–commutative black hole
Initially, the analysis begins by outlining the underlying framework for treating gravity through the lens of a non–commutative gauge theory. The core of this construction lies the $\mathrm{SO}(4,1)$ group, associated with the de Sitter symmetry, which serves as the foundational gauge group. As a starting point, the formalism is developed in a standard (3+1)(3+1)–dimensional Minkowski spacetime, still within a commutative setting. Within this background, the line element in spherical coordinates takes the familiar form:
@@
The metric’s angular sector is described by $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\varphi^{2}$, as one should expect. As mentioned before, the underlying symmetry is governed by the $\mathrm{SO}(4,1)$ group, whose algebra is generated by ten antisymmetric elements, labeled $\mathcal{M}_{ab}$ and satisfying the condition $\mathcal{M}_{ab}=-\mathcal{M}_{ba}$. Here, the indices $a$ and $b$ span the range 0,1,2,3,5{0,1,2,3,5}, with the subset 0,1,2,3{0,1,2,3} denoted by lowercase Latin indices from the beginning of the alphabet. These generators can be separated into two distinct classes: $\mathcal{M}_{ab}$ (with $a,b=0,1,2,3$), associated with local Lorentz rotations, and $P_{a}=\mathcal{M}_{a5}$, which fundamentally are generators of spacetime translations.
@@
[…]
@@
[…]
@@
Within this framework, the parameter $M$ identifies the mass of the black hole, while $\ell$ serves as the deformation parameter associated with the Kalb–Ramond sector Yang:2023wtu . The modified line element is written as $\mathrm{d}s^{2}=g^{(\Theta,\ell)}_{\mu\nu}(x,\Theta)\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}$, where the coordinates $x^{\mu}=(t,r,\theta,\varphi)$ describe a (3+1)(3+1)–dimensional spacetime endowed with non–commutative structure. Constructing this geometry involves determining the deformed tetrad fields $\hat{e}^{a}_{\mu}(x,\Theta)$, which arise through a contraction mechanism linking the non–commutative gauge group $\mathrm{SO}(4,1)$ to the Poincaré group $\mathrm{ISO}(3,1)$, following the Seiberg–Witten map formalism 5 ; 4 ; 3 . The resulting spacetime, modified by non–commutativity, is subject to the following structural constraints:
@@
$\omega^{ab}_{\mu\nu\rho}(x)=\frac{1}{4}\left\{\tilde{\omega}_{\nu},\partial_{\rho}\tilde{\omega}_{\mu}+R_{\rho\mu}\right\}^{ab},$ (13)
@@
+ $\quad\;+\;2\,\partial_{\nu}\,\tilde{\omega}_{\lambda}^{a\,b}\,\partial_{\rho}\partial_{\tau}\,e_{\mu}^{c}-\,2\,\partial_{\rho}\Bigl{(}\partial_{\tau}\,\tilde{\omega}_{\mu}^{a\,b}+R_{\tau\mu}^{a\,b}\Bigr{)}\,\partial_{\nu}\,e_{\lambda}^{c}\;-\;\Bigl{\{}\tilde{\omega}_{\nu},\,\bigl{(}\partial_{\rho}\,\tilde{\omega}_{\lambda}+R_{\rho\lambda}\bigr{)}\Bigr{\}}^{a\,b}\,\partial_{\tau}\,e_{\mu}^{c}$
@@
Here, the symbol $\ast$ stands for the conventional Moyal (star) product used in non–commutative geometry. To obtain the black hole solution, we adopt the Moyal twist $\partial_{r}\wedge\partial_{\theta}$, from which the corresponding matrix deformation is derived as ($\Theta=\Theta^{23}=-\Theta^{32}$)
$\Theta^{\mu\nu}=\left(\begin{matrix}0&0&0&0\\ 0&0&\Theta&0\\ 0&-\Theta&0&0\\ 0&0&0&0\end{matrix}\right),\qquad\mu,\nu=0,1,2,3.$ (24)
In other words, this prescription gives rise to the following commutation relation:
$r\,\overset{\star}{,}\,\theta\,]=i\Theta.$ (25)
Throughout the remainder of this analysis, we adopt natural units by setting $\hbar=c=G=1$. The corresponding line element can now be written explicitly as follows:
@@ III.1 Hawking temperature
Notice that by taking into account the temperature in terms of the mass is essential for probing whether a remnant mass ($M_{\text{rem}}$) emerges and for evaluating the black hole’s evaporation dynamics within the framework of the Stefan–Boltzmann law. While the detailed thermodynamic evolution—including aspects such as greybody factors, and absorption cross sections—is of considerable interest, these topics extend beyond the objectives of the present work. It is important to emphasize that, as the effective potential will be constructed through a perturbative approach (as discussed later), determining these quantities—particularly the last one—is far from straightforward. The complexity involved, therefore, makes the calculation highly nontrivial.
@@
$M^{1}_{\text{rem}}=\frac{i\sqrt{3}\Theta}{{4(1-\ell)}\sqrt{\ell+1}},$ (37)
$M^{2}_{\text{rem}}=\frac{i\sqrt{3}\Theta}{4(\ell-1)\sqrt{\ell+1}},$ (38)
or, to simplify the notation, one may also express it as:
$M^{(1,2)}_{\text{rem}}=\pm\frac{i\sqrt{3}\Theta}{4(\ell-1)\sqrt{\ell+1}}.$ (39)
@@ III.2 Heat capacity
$\begin{split}C^{(\Theta,\ell)}_{V}&=T\frac{\partial S}{\partial T}=-2\left(\pi r_{h}^{2}\right)-\frac{2\pi r^{2}\left(3\Theta^{2}+4(\ell+1)r_{h}^{2}\right)}{9\Theta^{2}+4(\ell+1)r_{h}^{2}}\\ &\approx\,-2\left(\pi r_{h}^{2}\right)+{3\pi(1-\ell)}\Theta^{2}.\end{split}$ (40)
@@ IV.1 Bosonic particle modes
$\mathcal{S}=\int p_{\mu}\,\mathrm{d}x^{\mu},$ (42)
where $p_{\mu}$ is the conjugate momentum.
When extracting the corresponding imaginary part, $\text{Im}\,\mathcal{S}$, it is the radial contribution that governs the result. The temporal term $p_{t}\,\mathrm{d}t=-\omega\,\mathrm{d}t$ remains purely real and, therefore, does not contribute to the tunneling amplitude. As such, only the radial integration yields a nonzero imaginary component, which enable us to evaluate the tunneling probability as follows
@@
Let us define the Hamiltonian as $H=M-\omega^{\prime}$, with $\omega^{\prime}$ denoting the instantaneous energy of the particle being emitted. According to Hamilton’s equations, this leads to the differential relation $\mathrm{d}H=-\mathrm{d}\omega^{\prime}$. Integrating over the entire energy range of the emitted particle, from 0 to the total energy $\omega$, the imaginary part of the action becomes:
@@
$\begin{split}\frac{\mathrm{d}r}{\mathrm{d}t}=&-h(r)+\sqrt{f(r)+h(r)^{2}}\\ &=\frac{1}{2}\left(\sqrt{\frac{\gamma(\Theta,\ell)\left(2r^{3}(2(\ell-1)M+r)+\Theta^{2}M(-3(\ell-1)M-2r)\right)}{r^{6}(2(\ell-1)M+r)^{2}}}\right.\\ &\left.-\sqrt{\frac{\gamma(\Theta,\ell)\left(2r^{2}(2(\ell-1)M+r)(2(\ell-1)M+\ell r)-\Theta^{2}(\ell-1)M(3(\ell-1)M+2r)\right)}{(\ell-1)r^{6}(2(\ell-1)M+r)^{2}}}\right)\\ &\approx\,1-\frac{\sqrt{\frac{2\ell M+\ell r-2M}{\ell-1}}}{\sqrt{r}}+\frac{(4\ell M^{2}-4M^{2}+Mr)\Theta^{2}}{2\,r^{3}(2\ell M-2M+r)}\\ &-\frac{\sqrt{\frac{2\ell M+\ell r-2M}{\ell-1}}\left(22\ell^{2}M^{3}+8\ell^{2}M^{2}r-44\ell M^{3}+3\ell M^{2}r+2\ell Mr^{2}+22M^{3}-11M^{2}r+2Mr^{2}\right)\Theta^{2}}{4r^{7/2}(2\ell M-2M+r)(2\ell M+\ell r-2M)},\end{split}$ (45)
@@
$\begin{split}&\text{Im}\,\mathcal{S}=\text{Im}\,\int_{0}^{\omega}-\mathrm{d}\omega^{\prime}\\ &\times\int_{r_{i}}^{r_{f}}\,\frac{\mathrm{d}r}{1-\frac{\sqrt{\frac{2\ell(M-\omega^{\prime})+\ell r-2(M-\omega^{\prime})}{\ell-1}}}{\sqrt{r}}+\frac{(4\ell(M-\omega^{\prime})^{2}-4(M-\omega^{\prime})^{2}+(M-\omega^{\prime})r)\Theta^{2}}{2\,r^{3}(2\ell(M-\omega^{\prime})-2(M-\omega^{\prime})+r)}+\Xi(\Theta,\ell,\omega^{\prime})},\end{split}$ (46)
@@
$\begin{split}\Gamma(\Theta,\ell,\omega)&\sim e^{-2\,\text{Im}\,S}=e^{-4\pi(1-\ell)^{2}\omega(2M-\omega)-\frac{\pi(16\ell-17)\big{[}\ln M-\ln(M-\omega)\big{]}\Theta^{2}}{8(1-\ell)^{2}}}\\ &{\approx\,\,8\pi\ell\omega e^{-4\pi\omega(2M-\omega)}(2M-\omega)+e^{-4\pi\omega(2M-\omega)}}\\ &{\Bigg{[}\frac{1}{4}\pi\ell e^{-4\pi\omega(2M-\omega)}\left(136\pi M\omega-68\pi\omega^{2}+9\right)(\ln M-\ln(M-\omega))}\\ &{+\frac{17}{8}\pi e^{-4\pi\omega(2M-\omega)}(\ln M-\ln(M-\omega))\Bigg{]}\Theta^{2}}.\end{split}$ (53)
@@
$\begin{split}n(\Theta,\ell,\omega)&=\frac{\Gamma(\Theta,\ell,\omega)}{1-\Gamma(\Theta,\ell,\omega)}=\frac{1}{\exp\left\{4\pi(1-\ell)^{2}\omega(2M-\omega)+\frac{\pi(16\ell-17)\big{[}\ln M-\ln(M-\omega)\big{]}\Theta^{2}}{8(1-\ell)^{2}}\right\}-1}\\ &{\approx\,\,\frac{8\pi\ell\omega e^{4\pi\omega(2M-\omega)}(2M-\omega)}{\left(e^{4\pi\omega(2M-\omega)}-1\right)^{2}}+\frac{1}{e^{4\pi\omega(2M-\omega)}-1}}\\ &{+\Bigg{[}\frac{17\pi e^{4\pi\omega(2M-\omega)}(\ln M-\ln(M-\omega))}{8\left(e^{4\pi\omega(2M-\omega)}-1\right)^{2}}}\\ &{+\frac{\pi e^{4\pi\omega(2M-\omega)}}{4\left(e^{4\pi\omega(2M-\omega)}-1\right)^{3}}\times\Bigg{(}-68\pi\omega^{2}e^{4\pi\omega(2M-\omega)}+136\pi M\omega e^{4\pi\omega(2M-\omega)}+136\pi M\omega}\\ &{+9e^{4\pi\omega(2M-\omega)}-68\pi\omega^{2}-9\Bigg{)}\times\Bigg{(}\ell(\ln M-\ln(M-\omega))\Bigg{)}\Bigg{]}\Theta^{2}}.\end{split}$ (54)
Fig. 6 presents the distribution of bosonic particle production, denoted by $n(\Theta,\ell,\omega)$111 It should be emphasized that, while the expansion in the second line of Eq. (54) is written only up to second order in $\Theta$ and first order in $\ell$, the plots in Fig. 6 were produced using the complete expression. This was done so that the effects of non–commutativity and Lorentz violation appear more clearly and can be more readily distinguished in the figures. . As observed, both the non–commutative parameter $\Theta$ and the Lorentz–violating parameter $\ell$ increase the particle creation density. However, it is worth emphasizing that the system exhibits a stronger response to variations in $\ell$ than to changes in $\Theta$.